OpenStudy (anonymous):

a hot air balloon is descending at a constant 6 m/s when you throw your iPhone horizontally with an initial 8 m/s. Find the magnitude and the direction of your phones velocity. Find the horizontal component of your phone's velocity just before it hits the ground?

OpenStudy (anonymous):

you have two velocities in different directions. This means you must use the component displacement equations 

OpenStudy (anonymous):



OpenStudy (anonymous):

using Pythagorean theorem, you get \[v_0^2=6^2+8^2\] \[v_0=\sqrt{36+64}=\sqrt{100}=10\] so this is th magnitude of the initial velocity of the phone

OpenStudy (anonymous):

Use the trig functions to find the direction... you know the the sides and are looking for theta

OpenStudy (anonymous):

lastly I believe there is missing information for the last question

OpenStudy (anonymous):

thank you